# Speculative Ensemble: Fast Large Language Model Ensemble via Speculation

### v1

@@ Abstract
Ensemble methods enhance Large Language Models (LLMs) by combining multiple models but suffer from high computational costs. In this paper, we introduce Speculative Ensemble, a novel framework that accelerates LLM ensembles without sacrificing performance, inspired by Speculative Decoding—where a small proposal model generates tokens sequentially, and a larger target model verifies them in parallel. Our approach builds on two key insights: (1) the verification distribution can be the ensemble distribution of both the proposal and target models, and (2) alternating each model as the proposer and verifier can further enhance efficiency. We generalize this method to ensembles with $n$ models and theoretically prove that SE is never slower than a standard ensemble, typically achieving faster speed. Extensive experiments demonstrate speed improvements of 1.11x–2.23x over standard ensemble techniques without compromising generation quality. Our code is available at https://github.com/Kamichanw/Speculative-Ensemble/.
Speculative Ensemble, Large Language Models, Inference Acceleration, Model Ensembles
@@ 1 Introduction
Recent advancements in Large Language Models (LLMs) have been substantial, with both closed-source and open-source frameworks pushing the boundaries of performance. Closed-source models, such as GPT-4 (Achiam et al., 2023) and the Gemini series (Team et al., 2023), have set high-performance benchmarks, while open-source models like Llama (Dubey et al., 2024), Deepseek (Liu et al., 2024), and the Qwen series (Bai et al., 2023) not only approach these benchmarks but also bring diversity in architecture, training data, and strategies. This variability has spurred interest in ensemble approaches that leverage the strengths of multiple models. Beyond traditional methods that average the probabilities of LLMs with similar capabilities, more sophisticated ensemble strategies have emerged. These include integrating three or more models to enhance performance (Yu et al., 2024) and using contrastive decoding (Li et al., 2023), which involves subtracting the logits of a smaller LLM from a larger one to reduce biases.
Despite the significant progress in ensemble techniques, a persistent challenge remains. Specifically, when combining multiple models for prediction, each LLM must sequentially compute the token distribution, leading to a considerable slowdown compared to the inference time of a single model. This raises the crucial question of whether ensemble speed can be increased without sacrificing the ensemble quality. To address this issue, we introduce Speculative Ensemble (SE), a novel ensemble approach inspired by Speculative Decoding (SD) (Leviathan et al., 2023).
@@
Figure 1: Comparison of (a) vanilla ensemble, (b) speculative decoding, and (c) speculative ensemble. In (b) and (c), each discrete blue block represents a probability calculated by one forward pass of $\mathcal{M}_{q}$, while the continuous green block indicates the joint distribution requires only one forward pass of $\mathcal{M}_{p}$.
@@
SD is a technique designed to accelerate LLM inference without sacrificing performance. As depicted in  Figure 1 (b), it uses a smaller but more efficient proposal model $\mathcal{M}_{q}$ to rapidly generate proposal tokens, which are then verified in parallel by a larger target model $\mathcal{M}_{p}$. The target model accepts a subset of these proposal tokens, enabling the generation of multiple tokens in a single forward pass, thus significantly accelerating inference. Moreover, by employing specific acceptance-rejection criteria, the generated tokens can be considered as samples drawn from the distribution of the target model, thus ensuring the generation quality. In this paper, we extend speculative decoding to model ensembles based on the following two observations.
First, SD allows not only sampling from the target model’s distribution, but also sampling from the ensemble distribution of the proposal model and target model. In vanilla SD, the target model’s distribution is directly employed for token verification and resampling, ensuring that the generated tokens align with the target model’s distribution. Similarly, we find that if the ensemble distribution is used for verification and resampling, as illustrated in  Figure 1 (c), the generated tokens will follow the ensemble distribution. Furthermore, this speculative-based ensemble significantly reduces the number of model invocations required. For instance, as shown in  Figure 1 (a), generating four tokens with the vanilla ensemble necessitates four invocations to both $\mathcal{M}_{q}$ and $\mathcal{M}_{p}$, while the speculative-based ensemble, in the optimal case, requires only four invocations to $\mathcal{M}_{q}$ and a single invocation to $\mathcal{M}_{p}$.
@@
Secondly, alternating each model as proposer and verifier can further accelerate the ensemble process. In standard SD, the proposer and verifier are fixed, with one model consistently serving as the proposer and the other as the verifier. However, we observe that in the ensemble setting, this static assignment is suboptimal, as it fails to fully leverage the bonus token. In SD, when all tokens from the proposal model are accepted by the target model, the target model will naturally generate an additional token, referred to as the bonus token. However, since the bonus token is drawn from the target model’s distribution rather than the ensemble distribution, it cannot be directly appended to the ensemble output. A naive solution might be to discard the bonus token or to re-query the proposal model and compute the ensemble distribution. Instead, we propose an alternative approach: treating the bonus token as a proposal from the target model, which is then verified by the proposer model. This insight leads to the development of a more efficient framework, the Alternate Proposal Framework, illustrated in Figure 2. Combined with the speculative-based ensemble, the alternate proposal framework forms the proposed SE. Moreover, we extend SE to the more general $n$-model ensemble case in Section 3.4.
@@
We establish the effectiveness of SE through both theoretical and experimental perspectives. Theoretically, we derive an expected improvement factor to quantify its acceleration and prove that SE is guaranteed to be at least as efficient as the standard ensemble, typically achieving greater speed. Additionally, in the weighted ensemble setting—the most common ensemble setting—we demonstrate that SE maintains a provable lower bound on the acceptance rate, ensuring consistently high efficiency. Experimentally, we conduct extensive experiments across various tasks, including code generation, mathematical reasoning, multi-task understanding, and text summarization. Our evaluation covers multiple LLM pairs, including Llama, Vicuna, and Qwen series, under both two-model and three-model configurations. The results show that SE consistently achieves the highest acceleration, with speedups of 1.34x–1.85x for weighted ensemble and 1.11x–2.23x for contrastive decoding.
@@
In summary, our key contributions are as follows: (1) We extend speculative decoding to the ensemble setting by refining its verification mechanism, introducing a speculative-based ensemble that significantly improves efficiency. (2) We incorporate an alternate proposal framework into the speculative-based ensemble to get finally Speculative Ensemble, further boosting inference speed. (3) Through extensive theoretical analysis and experimental evaluation, we demonstrate that our method achieves substantial acceleration while maintaining a lower bound, ensuring it never underperforms compared to standard ensembles.
@@ 2 Related Work
Large Language Model Ensemble. LLM ensembles combine multiple pretrained models to enhance output quality and robustness. Existing approaches can be categorized based on the application stage: pre-inference, post-inference, and during-inference (Lu et al., 2024a).
Pre-inference ensembles use a router to direct incoming queries to the most suitable LLM (Shnitzer et al., 2023; Lu et al., 2023, 2024b; Ding et al., 2024). Post-inference ensembles operate at the sequence level, selecting the optimal output from multiple LLM-generated sequences (Chen et al., 2023b; Lee et al., 2023; Jiang et al., 2023). We focus on during-inference ensembles, where multiple LLMs contribute at the token level by leveraging probability distributions or logits to predict the next token. Methods in this category include weighted sums of probability distributions or logits (Li et al., 2024a), contrastive decoding (Li et al., 2023), and learnable ensemble techniques (Furkan Tekin et al., 2024). The proposed SE can accelerate any probability or logits-level ensemble, demonstrating broad potential.
Speculative Decoding. Speculative decoding (Leviathan et al., 2023; Chen et al., 2023a) can be categorized into two main areas: proposal model design and verification design. In the first category, proposal models are designed to generate tokens that are more likely to be accepted by the verifier. This includes independent proposal models, such as distillation-based method (Zhou et al., 2024) and target-informed models that incorporate information of verifier (Zhang et al., 2024; Elhoushi et al., 2024; Monea et al., 2023; Yi et al., 2024; Monea et al., 2023; Li et al., 2024b).
@@
The most related work is Speculative Contrastive Decoding (SCD) (Yuan et al., 2024), which improves output quality by using contrastive decoding for rejected tokens. However, SE applies to any ensemble function allowing application to all tokens instead of only rejected ones.
3 Speculative Ensemble
@@ 3.1 Speculative Decoding
Speculative decoding (SD) is a technique designed to speed up inference while maintaining the quality of generated outputs. It involves two phases: the proposal phase and the verification phase. During the proposal phase, a lightweight proposal model sequentially generates proposal tokens. In the verification phase, a larger target model verifies these tokens in parallel. Furthermore, by incorporating appropriate acceptance-rejection criteria, the technique ensures that the generated tokens align precisely with the target model’s distribution, thus maintaining high-quality results.
@@
3.2 Speculative-based Ensemble
As discussed above, vanilla SD can only accelerate the inference of a single model. In this subsection, we will introduce how to apply SD to scenarios involving an arbitrary ensemble of two models. Specifically, let $q_{i}(x)$ and $p_{i}(x)$ denote the distributions of token $x_{i}$ given by the proposal model and the target model, respectively, and let $l^{q}_{i}$ and $l^{p}_{i}$ be the corresponding logits. Then, the ensemble distribution $r_{i}(x)$ can be expressed as
$r_{i}(x)=\mathcal{E}(q_{i}(x),p_{i}(x))\ \ \text{or}\ \ \mathcal{E^{\prime}}(l^{q}_{i},l^{p}_{i}),$ (3)
where $\mathcal{E}(\cdot)$ represents the ensemble function at the probability level, while $\mathcal{E}^{\prime}(\cdot)$ is at logits level. For example, the common weighted ensemble form that uses probability for weighted summation can be expressed as
$r(x)=\mathcal{E}(q_{i}(x),p_{i}(x))=\lambda q(x)+(1-\lambda)p(x),$ (4)
@@ 3.2 Speculative-based Ensemble
$r(x)=\mathcal{E^{\prime}}(l^{q}_{i},l^{p}_{i})=\text{Softmax}(l^{p}_{i}-\mu l^{q}_{i}).$ (5)
We note that by making slight modifications to the vanilla SD, the generated tokens can align with the ensemble distribution. Specifically, before verification, we first compute the ensemble distribution $r(x)$ and update the verification formula in  Equation 2 as follows:
@@
We theoretically prove the correctness of speculative-based ensemble, that is, the tokens generated by the above sampling process precisely align with the ensemble distribution, with an acceptance rate $\alpha$ of
@@
The proof of Theorem 3.1 is in Section A.2. Theorem 3.1 provides the speed improvement factor when $\alpha$ is known. However, in most cases, $\alpha$ is unknown and requires extensive experiments to estimate. Nevertheless, we find that in the weighted ensemble scenario, which is the most common ensemble scenario, $\alpha$ has a lower bound.
@@
If $\mathcal{E}(p,q)=\lambda q(x)+(1-\lambda)p(x)$ and $q(x)$ is the proposal model. Then $\alpha$ has a lower bound of $\lambda$.
@@
Assume that $\mathcal{E}(p,q)=\lambda q(x)+(1-\lambda)p(x)$, and that $\mathcal{M}_{q}$ and $\mathcal{M}_{p}$ have comparable parameters. Then, by selecting an appropriate proposal model, $\alpha$ has a lower bound of at least 0.5.
@@
By utilizing the lower bound property, we demonstrate that the proposed speculative-based decoding method is guaranteed to be no slower than the vanilla ensemble approach in the weighted ensemble scenario, and it is typically faster.
@@
Assume that $\mathcal{E}(p,q)=\lambda q(x)+(1-\lambda)p(x)$, then if $\lambda>\frac{c}{1+c}$, there exists a value of $\gamma$ that enhances the inference speed.
@@
Assume that $\mathcal{E}(p,q)=\lambda q(x)+(1-\lambda)p(x)$. Then, for any two models, there exists a value of $\gamma$ such that the speed of the speculative-based ensemble is not slower than the vanilla ensemble, and it is almost always faster.
@@
Interpretable quality-speed tradeoff in vanilla SD. In SD, some studies focus on relaxing the acceptance criteria for a higher acceptance rate to achieve faster inference, such as lossy SD (Zhou et al., 2024) and typical acceptance (Cai et al., 2024). However, these methods often lack interpretability, that is, we do not know which distribution the generated tokens will follow. We find that the speculative-based ensemble can naturally be an interpretable strategy for adjusting the quality-speed tradeoff in vanilla SD.
Specifically, we apply a weighted ensemble using the proposal and target models in SD. In this setup, when $\lambda=0$, the ensemble distribution aligns exactly with the target distribution, reducing the method to standard SD. For $\lambda>0$, the acceptance rate is guaranteed to have a lower bound and greater than that of vanilla SD, as demonstrated in the proof of Theorem 3.2, leading to greater acceleration. However, incorporating less precise information from a smaller model can introduce some performance degradation. This tradeoff provides a mechanism to balance quality and speed in speculative decoding.
In contrast to existing approaches, this method improves interpretability. This is because we know the distribution of the generated tokens after relaxation, i.e. ensemble distribution defined in Equation 4. This allows us to design proposal models that accelerate inference without compromising performance, and potentially even enhance the model’s capabilities in specific areas. For instance, some research suggests that ensembling a smaller model appropriately can improve safety (Wang et al., 2024; Li et al., 2024a). The experimental results are shown in Section C.1.

### v2

# Fast Large Language Model Collaborative Decoding via Speculation

@@ Abstract
Large Language Model (LLM) collaborative decoding techniques improve output quality by combining the outputs of multiple models at each generation step, but they incur high computational costs. In this paper, we introduce Collaborative decoding via Speculation (CoS), a novel framework that accelerates collaborative decoding without compromising performance. Inspired by Speculative Decoding—where a small proposal model generates tokens sequentially, and a larger target model verifies them in parallel, our approach builds on two key insights: (1) the verification distribution can be the combined distribution of both the proposal and target models, and (2) alternating each model as the proposer and verifier can further enhance efficiency. We generalize this method to collaboration among $n$ models and theoretically prove that CoS is never slower than standard collaborative decoding, typically achieving faster speed. Extensive experiments demonstrate CoS is 1.11x–2.23x faster than standard collaborative decoding without compromising generation quality. Our code is available at https://github.com/Kamichanw/CoS/.
Collaborative Decoding, Speculative Decoding, Inference Acceleration, Large Language Models
@@ 1 Introduction
Recently, large language models (LLMs) have demonstrated impressive performance across a wide range of tasks. Beyond advances in individual models—such as architectural innovations and training techniques—there is increasing interest in collaborative approaches involving multiple LLMs (Lu et al., 2024; Chen et al., 2025). A key class of these methods combines information from multiple models (e.g., probability distributions or logits) during token generation to improve next-token prediction and capitalize on the complementary strengths of different models. For instance, ensembling methods (Yu et al., 2024; Huang et al., 2024; Yao et al., 2024) average prediction distributions from multiple models; contrastive decoding (Li et al., 2023; O’Brien & Lewis, 2023) improves generation quality and reduces hallucinations by subtracting the outputs of a smaller model from a larger one; and decoding-time realignment (Liu et al., 2024; Shi et al., 2024) jointly uses an aligned and an unaligned model during decoding to enable flexible control over alignment. In this paper, we refer to such approaches collectively as collaborative decoding.
@@
Figure 1: Comparison of (a) vanilla collaborative decoding, (b) speculative decoding, and (c) collaborative decoding via speculation. In (b) and (c), each discrete blue block represents a probability calculated by one forward pass of $\mathcal{M}_{q}$, while the continuous green block indicates the joint distribution requires only one forward pass of $\mathcal{M}_{p}$.
Despite the significant progress in collaborative decoding, a key challenge persists: combining outputs from multiple models requires each model to perform a separate forward pass, which substantially slows down inference compared to using a single model. This raises a crucial question: can we speed up collaborative decoding without compromising quality? To address this, we propose Collaborative decoding via Speculation (CoS), a novel framework that accelerates any form of collaborative decoding while maintaining output quality. The core idea of CoS comes from Speculative Decoding (SD) (Xia et al., 2023; Leviathan et al., 2023).
@@
Speculative Decoding is a technique designed to accelerate LLM inference without sacrificing performance. As depicted in  Figure 1 (b), it uses a smaller but more efficient proposal model $\mathcal{M}_{q}$ to rapidly generate proposal tokens, which are then verified in parallel by a larger target model $\mathcal{M}_{p}$. The target model accepts a subset of these proposal tokens, enabling the generation of multiple tokens in a single forward pass, thus significantly accelerating inference. Moreover, by employing specific acceptance-rejection criteria, the generated tokens can be considered as samples drawn from the distribution of the target model, thus ensuring the generation quality. In this paper, we extend speculative decoding to LLM collaborative decoding based on the following two observations.
(a) Standard Collaborative Decoding
(b) Alternate Proposal Framework
@@
First, SD allows not only sampling from the target model’s distribution, but also sampling from any combined distribution of the proposal model and target model. In vanilla SD, the target model’s distribution is directly employed for token verification and resampling, ensuring that the generated tokens align with the target model’s distribution. Similarly, we find that if the combined distribution is used for verification and resampling, as illustrated in Figure 1 (c), the generated tokens will follow the combined distribution. We refer to this generalization of SD to collaborative decoding as Naive-CoS. Naive-CoS significantly reduces the number of model invocations required. For instance, as shown in Figure 1 (a), generating four tokens with the vanilla collaborative decoding necessitates four invocations to both $\mathcal{M}_{q}$ and $\mathcal{M}_{p}$, while Naive-CoS, in the optimal case, requires only four invocations to $\mathcal{M}_{q}$ and a single invocation to $\mathcal{M}_{p}$.
@@
Secondly, alternating each model as proposer and verifier can further accelerate the collaboration process. In standard SD, the proposer and verifier are fixed, with one model consistently serving as the proposer and the other as the verifier. However, we observe that in the collaborative decoding setting, this static assignment is suboptimal, as it fails to fully leverage the bonus token. In SD, when all tokens from the proposal model are accepted by the target model, the target model will naturally generate an additional token, referred to as the bonus token. However, since the bonus token is drawn from the target model’s distribution rather than the combined distribution, it cannot be directly appended to the output of collaborative decoding. A naive solution might be to discard the bonus token or to re-query the proposal model and compute the combined distribution. Instead, we propose a more efficient approach: treating the bonus token as a proposal from the target model, which is then verified by the proposer model. This insight leads to the Alternate Proposal Framework, illustrated in Figure 2. Combined with the Naive-CoS, the alternate proposal framework forms the proposed CoS. We further extend CoS to the general case of $n$-model collaboration in Section 3.4.
As shown in Figure 2(a), in standard collaborative decoding, each model invocation can generate only one probability distribution, so generating $n$ tokens requires $2n$ model invocations. With the alternate proposal framework (Figure Figure 2(b)), each invocation can generate two distributions in the optimal case, thereby doubling the generation efficiency.
@@
We establish the effectiveness of CoS through both theoretical and experimental perspectives. Theoretically, we derive an expected improvement factor to quantify its acceleration and prove that CoS is guaranteed to be at least as efficient as the standard collaborative decoding, typically achieving greater speed. Additionally, in the weighted ensemble setting—the most common collaborative decoding setting—we demonstrate that CoS maintains a provable lower bound on the acceptance rate, ensuring consistently high efficiency. Experimentally, we conduct extensive experiments across various tasks, including code generation, mathematical reasoning, multi-task understanding, and text summarization. Our evaluation covers multiple LLM pairs, including Llama, Vicuna, and Qwen series, under both two-model and three-model configurations. The results show that CoS consistently achieves the highest acceleration, with speedups of 1.34x–1.85x for weighted ensemble and 1.11x–2.23x for contrastive decoding.
@@
In summary, our key contributions are as follows: (1) We extend speculative decoding to the collaborative decoding setting by refining its verification mechanism, introducing a Naive-CoS that significantly improves efficiency. (2) We incorporate an alternate proposal framework into the Naive-CoS to get the final CoS, further boosting inference speed. (3) Through extensive theoretical analysis and experimental evaluation, we demonstrate that our method achieves substantial acceleration while maintaining a lower bound, ensuring it never underperforms compared to standard collaborative decoding.
@@ 2 Related Work
LLM Collaborative Decoding. In this paper, LLM collaborative decoding refers to the integration of information from multiple LLMs when generating the next token. This typically involves combining the output probabilities or logits to compute a final probability distribution. This method can be used to improve overall performance, especially in solving complex tasks(Han et al., 2025, 2024). A common approach is model ensembling, which averages or applies weighted averaging to the probability distributions of multiple LLMs to derive the final sampling distribution (Yu et al., 2024; Huang et al., 2024; Yao et al., 2024). Studies have shown that this technique can enhance both performance and safety (Li et al., 2024a). Another method, contrastive decoding, is based on the observation that smaller models tend to produce noisier outputs. By subtracting the logits of a smaller model from those of a larger one, a cleaner and more reliable set of logits can be obtained, resulting in higher-quality outputs (Li et al., 2023; O’Brien & Lewis, 2023). Finally, decoding-time realignment enables flexible alignment with human preferences during decoding by linearly combining the logits of a human-aligned and an unaligned model, thereby balancing performance with alignment objectives (Liu et al., 2024; Shi et al., 2024).
Our proposed method, CoS, represents an orthogonal approach to existing collaborative decoding techniques. It is designed to substantially improve inference speed while preserving the benefits of collaboration. Importantly, CoS is not restricted to accelerating the three methods discussed above; its generality allows it to enhance the efficiency of any collaborative decoding approach, including those yet to be developed.
Speculative Decoding. Speculative decoding (Xia et al., 2023; Leviathan et al., 2023; Chen et al., 2023) can be categorized into two main areas: proposal model design and verification design. In the first category, proposal models are designed to generate tokens that are more likely to be accepted by the verifier. This includes independent proposal models, such as distillation-based method (Zhou et al., 2024) and target-informed models that incorporate information of verifier (Zhang et al., 2024; Elhoushi et al., 2024; Monea et al., 2023; Yi et al., 2024; Monea et al., 2023; Li et al., 2024b; Sun et al., 2024b).
@@
The most closely related work is Speculative Contrastive Decoding (SCD) (Yuan et al., 2024), which combines outputs from both large and small models during the verification phase to form a contrastive decoding distribution. Operationally, SCD can be seen as a special case of Naive-CoS in the contrastive decoding setting. The method proposed in this paper differs from SCD in three key ways: (1) CoS is more broadly applicable and can accelerate any collaborative decoding approach; (2) it introduces an alternative proposal framework that further improves decoding speed; and (3) it provides a complete theoretical analysis, ensuring inference efficiency comparable to standard collaborative decoding methods.
3 Collaborative Decoding via Speculation
@@ 3.1 Speculative Decoding
Unlike other acceleration methods (Sun et al., 2024a), speculative decoding (SD) is a technique designed to speed up inference while maintaining the quality of generated outputs. It involves two phases: the proposal phase and the verification phase. During the proposal phase, a lightweight proposal model sequentially generates proposal tokens. In the verification phase, a larger target model verifies these tokens in parallel. Furthermore, by incorporating appropriate acceptance-rejection criteria, the technique ensures that the generated tokens align precisely with the target model’s distribution, thus maintaining high-quality results.
@@
3.2 Naive-CoS
As discussed above, vanilla SD can only accelerate the inference of a single model. In this subsection, we will introduce how to apply SD to scenarios involving an arbitrary combination of two models. Specifically, let $q_{i}(x)$ and $p_{i}(x)$ denote the distributions of token $x_{i}$ given by the proposal model and the target model, respectively, and let $l^{q}_{i}$ and $l^{p}_{i}$ be the corresponding logits. Then, the combined distribution $r_{i}(x)$ can be expressed as
$r_{i}(x)=\mathcal{C}(q_{i}(x),p_{i}(x))\ \ \text{or}\ \ \mathcal{C^{\prime}}(l^{q}_{i},l^{p}_{i}),$ (3)
where $\mathcal{C}(\cdot)$ represents the combination function at the probability level, while $\mathcal{C}^{\prime}(\cdot)$ is at logits level. For example, the common weighted ensemble that uses probability for weighted summation can be expressed as
$r_{i}(x)=\mathcal{C}(q_{i}(x),p_{i}(x))=\lambda q_{i}(x)+(1-\lambda)p_{i}(x),$ (4)
@@ 3.2 Naive-CoS
$r_{i}(x)=\mathcal{C^{\prime}}(l^{q}_{i},l^{p}_{i})=\text{Softmax}(l^{p}_{i}-\mu l^{q}_{i}).$ (5)
We note that by making slight modifications to the vanilla SD, the generated tokens can align with the combined distribution. Specifically, before verification, we first compute the combined distribution $r_{i}(x)$ and update the verification formula in  Equation 2 as follows:
@@
We theoretically prove the correctness of Naive-CoS, that is, the tokens generated by the above sampling process precisely align with the combined distribution, with an acceptance rate $\alpha$ of
@@
The proof of Theorem 3.1 is in Section A.2. Theorem 3.1 provides the speed improvement factor when $\alpha$ is known. However, in most cases, $\alpha$ is unknown and requires extensive experiments to estimate. Nevertheless, we find that in the weighted ensemble scenario, which is the most common collaborative decoding, $\alpha$ has a lower bound.
@@
If $\mathcal{C}(p,q)=\lambda q(x)+(1-\lambda)p(x)$ and $q(x)$ is the proposal model. Then $\alpha$ has a lower bound of $\lambda$.
@@
Assume that $\mathcal{C}(p,q)=\lambda q(x)+(1-\lambda)p(x)$, and that $\mathcal{M}_{q}$ and $\mathcal{M}_{p}$ have comparable parameters. Then, by selecting an appropriate proposal model, $\alpha$ has a lower bound of at least 0.5.
@@
By utilizing the lower bound property, we demonstrate that the proposed Naive-CoS is guaranteed to be no slower than the weighted ensemble approach, and it is typically faster.
@@
Assume that $\mathcal{C}(p,q)=\lambda q(x)+(1-\lambda)p(x)$, then if $\lambda>\frac{c}{1+c}$, there exists a value of $\gamma$ that enhances the inference speed.
@@
Assume that $\mathcal{C}(p,q)=\lambda q(x)+(1-\lambda)p(x)$. Then, for any $\lambda$, there exists a value of $\gamma$ such that the speed of the Naive-CoS is not slower than the vanilla weighted ensemble, and it is almost always faster.
@@
Interpretable quality-speed tradeoff in vanilla SD. In SD, some studies focus on relaxing the acceptance criteria for a higher acceptance rate to achieve faster inference, such as lossy SD (Zhou et al., 2024) and typical acceptance (Cai et al., 2024). However, these methods often lack interpretability, that is, we do not know which distribution the generated tokens will follow. We find that Naive-CoS can naturally be an interpretable strategy for adjusting the quality-speed tradeoff in vanilla SD.
Specifically, we apply a weighted ensemble using the proposal and target models in SD. In this setup, when $\lambda=0$, the combined distribution aligns exactly with the target distribution, reducing the method to standard SD. For $\lambda>0$, the acceptance rate is guaranteed to have a lower bound and greater than that of vanilla SD, as demonstrated in the proof of Theorem 3.2, leading to greater acceleration. However, incorporating less precise information from a smaller model can introduce some performance degradation. This tradeoff provides a mechanism to balance quality and speed in speculative decoding.
In contrast to existing approaches, this method improves interpretability. This is because we know the distribution of the generated tokens after relaxation, i.e. combined distribution defined in Equation 4. This allows us to design proposal models that accelerate inference without compromising performance, and potentially even enhance the model’s capabilities in specific areas. For instance, some research suggests that ensembling a smaller model appropriately can improve safety (Wang et al., 2024; Li et al., 2024a). The experimental results are shown in Section C.1.